OpenStudy (anonymous):

let f(x)=1-2x^2 and g(x) = x+1 compute the following (g o f)(x) similar to the one I just did but backwords. ???

OpenStudy (anonymous):

always start this way \[g\circ f(x)=g(f(x))\] in other words get rid of the circle

OpenStudy (anonymous):

does it change things when they are in a different order?

OpenStudy (anonymous):

ok did that

OpenStudy (anonymous):

then replace the "inside function" by what you actually have \[g(1-2x^2)\] in this case

OpenStudy (anonymous):

oh yes it matter for sure!

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok got that

OpenStudy (anonymous):

now g says to take your input and add 1, so your answer is \[g(1-2x^2)=1-2x^2+1=2-2x^2\]

OpenStudy (anonymous):

think of it this way \[f(\heartsuit)=1-2\heartsuit^2\] \[g(\spadesuit)=\spadesuit +1\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

try not to think of the 'x" the variable is unimportant

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so if I have (g o f)(-1) I do the same steps as above ?

OpenStudy (anonymous):

\[f(\triangle)=1-2\triangle ^2\] now you can fill in the triangle with whatever you want

OpenStudy (anonymous):

oh no that means the inverse of the composition, that is harder

OpenStudy (anonymous):

oh great

OpenStudy (anonymous):

oh wait, is that simply \[g\circ f(-1)\]?

OpenStudy (anonymous):

it is not harder sorry, it is easier

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[g\circ f(-1)\] means 1: find f(-1) 2: find g of the result

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[f(-1)=1-2(-1)^2=1-2=-1\] and \[g(-1)=-1+1=0\] so \[g\circ f(-1)=0\]

OpenStudy (anonymous):

ok that is what i got

OpenStudy (anonymous):

To know why does it have to be (g o f) here's a picture of the thing.  So you'll have to map f first to g, then from g to h. Hence, if we're not using the composite function yet, it'll be this. f(1)=2 Then g(2)=4 Hence, we can see that we have to map the first function first. Take note that 2=f(1). Therefore, g[f(1)]=4 Hence, we write the function as either gf(1)=4 or (g o f)(1)=4.